A class has 29 students, of which 5 are seniors. How many committees of size 5 can be selected if at least one member of each committee must be a senior?

Answer by stanbon(75887)

A class has 29 students, of which 5 are seniors. How many committees of size 5 can be selected if at least one member of each committee must be a senior?
------------------------------------------------
P(no senior in group of 5) = 24C5/29C5 = 42504/118755 = 0.3579
----
P(at least one senior) = 1 - P(no senior)
= 1 - 0.3579
= 0.6421...
Cheers,
Stan H.